Which of the following is NOT an assumption of the traditional economic order quantity (EOQ) model?

a. Holding and ordering costs are stable and known.

b. Demand is constant and known.

c. Supply lead time is constant and known.

d. Quantity discounts are possible.

Kim’s Nail Salon uses a weighted moving average method to forecast demand. She assigns a weight of 5 to the previous month’s demand, 3 to demand two months ago, and 2 to demand three months ago. If demand was 800 customers in April, 900 customers in May, 1,200 customers in June, and 2,000 customers in July, what should her forecast for August be?

a. 1,540

b. 910

c. 1,400

d. 1,210

Statewide Insurers receives 600 applications per month. Currently, there are 60 applications in the Reception area and 140 applications with the Underwriting team. How long does Receiving take to process an application? Assume that there are 20 working days per month.

a. 2 days

b. 3 days

c. 7 days

d. 10 days

Based on experience, the manager at the Giraffe Hotel decides to accept a reservation for a party of 12 guests knowing that his hotel is sold out. He is accepting the reservation because:

a. he can cancel the reservations of the 12 lowest-paying guests and get more revenue from the new guests

b. he can make arrangements with the neighboring hotel to accept the 12 guests

c. he is using overbooking techniques and has determined that he can safely accept the 12 guests

d. he knows he will have 6 cancellations and 6 no-shows among the current reservations

A manufacturing firm finds a location using the center of gravity method but rejects that location because they are unable to efficiently transport the finished goods to market. The firm’s action illustrates the link between __________ and location.

a. proximity

b. clustering

c. infrastructure

d. globalization
Answer #1

Q1) The following is not an assumption of traditional economic order quantity (EOQ) model
d. Quantity discounts are possible

Q2)
Forecast for August = (5*2000+1200*3+900*2)/(5+3+2) = 1540

Ans. a) 1540

Q3)
Time receiving takes per application=(140+60)/20 = 10 days

Ans d. 10 days

Q4)
c. He is using overbooking techniques and has determined that he can safely accept the 12 guests

Q5)
It illustrates the link between
a proximity